f(x)=(x-6)/x^3 +5x
Answer by Fombitz(32388)


What's your question?
Also, use parentheses to distinguish what's in the denominator.
Is is f(x)=(x-6)/(x^3)+5x (red line) or is is f(x)=(x-6)/(x^3+5x) (green line)??
Please repost.